# On graphs with large third eigenvalue

### v3

@@ 5 Concluding remarks
Regarding Theorem 4.6, although we were unable to establish explicit bounds on the penultimate root of the characteristic polynomial of the vertex multiplications of $T$, we suspect that with further simplifications or insights, such bounds could be derived. For example, if we vertex multiply $T$ by $[x_{1},\dots,x_{11}]$, we see that $x_{11}$ does not affect the non-zero eigenvalues of $T$, thus we may assume that $x_{11}=0$. Note that the undirected, simple graph corresponding to $T$ has automorphism group isomorphic to $V_{4}\cong C_{2}\times C_{2}$ which simplifies matters slightly.
@@ 6 Acknowledgements
We pursued this research as part of the Summer Research in Mathematics (SRIM) programme organised by the Department of Pure Mathematics and Mathematical Sciences at the University of Cambridge. We were supervised by Jan Petr and are very grateful for the meet-ups, advice and insight which we found incredibly helpful.
@@
The characteristic polynomial of the vertex multiplication $\tilde{T}$ of $T$ by $[x_{1},x_{2},\dots,{x}_{10},x_{11}]$ is given by:
$\chi_{\tilde{T}}(t)=t^{11}$displaystyle-- $(x_{1}x_{5}+x_{1}x_{6}+x_{2}x_{5}+x_{1}x_{7}+x_{2}x_{6}+x_{3}x_{5}+x_{1}x_{8}+x_{2}x_{7}+x_{3}x_{6}+x_{4}x_{5}$

### v4

@@ 5 Concluding remarks
Regarding Theorem 4.6, although we were unable to establish explicit bounds on the penultimate root of the characteristic polynomial of the vertex multiplications of $T$, we suspect that with further simplifications or insights, such bounds could be derived. For example, if we vertex multiply $T$ by $x_{1},\dots,x_{11}]$, we see that $x_{11}$ does not affect the non-zero eigenvalues of $T$, thus we may assume that $x_{11}=0$. Note that the undirected, simple graph corresponding to $T$ has automorphism group isomorphic to $V_{4}\cong C_{2}\times C_{2}$ which simplifies matters slightly.
@@ 6 Acknowledgements
We pursued this research as part of the Summer Research in Mathematics (SRIM) programme organised by the Department of Pure Mathematics and Mathematical Sciences at the University of Cambridge. We were supervised by Jan Petr and are very grateful for the meet-ups, advice and insight which we found incredibly helpful. We would also like to thank the anonymous reviewers for a careful reading, valuable comments and useful suggestions.
@@
The characteristic polynomial of the vertex multiplication $\tilde{T}$ of $T$ by $x_{1},x_{2},\dots,{x}_{10},x_{11}]$ is given by:
$\chi_{\tilde{T}}(t)=t^{11}$ - $(x_{1}x_{5}+x_{1}x_{6}+x_{2}x_{5}+x_{1}x_{7}+x_{2}x_{6}+x_{3}x_{5}+x_{1}x_{8}+x_{2}x_{7}+x_{3}x_{6}+x_{4}x_{5}$